What is the concentration of a solution with a volume 3.3 mL that contains 12 grams of ammonium sulfite?


1
Expert's answer
2022-04-06T04:14:04-0400

Concentration= moles/ litre


Moles= 12/116= 0.3044moles


Molarity= 0.3M


For a solution of 3ml


(3.3/ 1000)×0.3= 0.001M